If the voltage of Zn-H+
cell is
0.542 V at 25 0 C, What is the
concentration of the [H+
]?
when : [Zn 2+] = 0.01 M and [H2
] =
1.0 M
1
Expert's answer
2021-10-08T02:08:49-0400

"E^o_{cell} = E^o_{reduction} (H^+\/H_2 ) - E^o_{reduction} (Zn^{2+} \/Zn)"

"= 0.00 V - (-0.76 V) \\

= 0.76 V \\

E_{cell} = 0.542 \\"

the overall equation is

Zn(s) + 2H+(aq) → Zn2+(aq) + H2(g)

so n = 2, and

So nernst equation :

"0.542 = 0.76 - (\frac{0.059}{2})logQ \\

-0.218 = -0.0295logQ \\

7.3898 = log Q \\

Q= 10^7.3898 = 24535787.41 = 2.45 \times 10^7"

where Q = ([Zn+2 ] [H2] )/[H+]2

"2.45 \times 10^7 = \frac{0.01 \times 0.1}{[H^+]^2}"

Now we solve for [H+]

"[H+] = \sqrt{ 0.408 \times 10^{-10}}= 0.638 \times 10^{-5} \\

= 6.38 \times 10^{-6} \;M"